1. find three consecutive even integers such that one fourth the sum of the first and the third is equal to 22 less than the second.
Thank you tutor!..:)

Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

find three consecutive even integers such that one fourth the sum of the first and the third is equal to 22 less than the second.
Thankyou tutor!..:)
Let x, x+2 and x+4 be the three consecutive even integers.
Then we have:
1/4*(x + (x+4)) = (x+2) - 22
Solve the above equation for x and then calculate x+2 and x+4.

Answer by richwmiller(17219)

(1/4)(n+n+4)=n+2-22
1/4*(2n+4)=n-20
n/2+1=n-20
n/2=n-21
n=2n-42
42=n
42,44,46
check
1/4*(42+46)=44-22
22=22
ok